OpenStudy (anonymous):

Solve the following system of equations. 9x + 5y = -132 4x + y = -33

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

You can use the substitution method. Take the second equation and solve it for the value of y. Then substitute that value in for y in the first equation so that you can solve it for x. Once you know what x is, substitute the value of x into the second equation to find y. 4x + y = -33 y = -33 - 4x 9x + 5(-33 - 4x) = -132 9x - 165 - 20x = -132 -11x = 33 x = -3 <--- 4(-3) + y = -33 -12 + y = -33 y = ? Hope this helps! :)